We had two problems during the night.
1. At the begging of the night, around UT 22:48, the telescope was slewing to observe the first target of the night. The target was the standard field E6-A at RA=14h45m DEC=-45d15m. The telescope started slewing, but half of the way the emergency brakes activated. The operator and technicians went upstairs to verify the origin of the problem and run several tests. At UT 01:00 we were able to restart the observations.
2.At the end of the night, around UT 9:01, the shutter didn't respond and the exposure was never taken. Alberto noticed that the vacuum of the camera was gone, I stoped the queue and click safe mode. The system became unresponsive and we couldn't turn off the vsub.
Total Times | Time Observed | Time Engineering | Time lost technical | Time lost weather | Time lost other | Total Program Time |
---|---|---|---|---|---|---|
2015-05-24 | 8.0 | 0.0 | 3.0 | 0.0 | 0.0 | 11 |
Links
[1] mailto:marla.geha@yale.edu
[2] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/2380/delete?destination=print/node/7811
[3] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/2380/edit?destination=print/node/7811
[4] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/add/node/7811?destination=print/node/7811
The first problem took about 1 hour and 10 min to be solved. The problem was solved and we were able to start observing.
The second problem was also solved and the system was restarted around 10:20 UT.